roy's welding supplies common stock sells for $42 a share in Paisan annual dividend that increases by 4.5% annually. The market rate of return on the stock is 9.20%, what is the amount of the last dividend paid?
Answer #1

Ans $ 1.89

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
42 = D1 / (9.20%- 4.5%)
D1 = 42 * (9.20%- 4.5%)
D1 = 1.974
D1 = D0* (1 + g)
1.974 = D0* (1 + 4.5%)
D0 = 1.974 / 1.045
D0 = 1.89
Answer #2

SOLUTION :


Current price, P0 = 42 ($)

Rate of return, r = 9.2 % = 0.092

Dividend growth rate, g = 4.5 % = 0.045


Let $ D0 be the last annual dividend paid.


Now,


D0 (1 + g) = (r - g) * P0

=> D0 * (1 + 0.045) = (0.092 - 0.045) * 42

=> 1.045 D0 = 1.974

=> D0 = 1.974/1.045

=> D0 = 1.89 ($) 


So, amount of the last annual dividend paid = $1.89 (ANSWER).



answered by: Tulsiram Garg